Given the following information, calculate the density in g/mL of an irregular solid.
Mass of weighing vessel 1.005g
Mass of solid + weighing vessel 9.441g
Volume of liquid in graduated cylinder 3.45 mL
Volume of liquid in graduated cylinder +
volume of solid
5.15 mL
Answer #1

Mass of weighing vessel      = 1.005g

Mass of solid + weighing vessel   = 9.441g

mass of solid    = (Mass of solid + weighing vessel)-Mass of weighing vessel

                         = 9.441-1.005   = 8.436g

Volume of liquid in graduated cylinder    = 3.45 mL

Volume of liquid in graduated cylinder +volume of solid       =   5.15 mL

volume of solid =( Volume of liquid in graduated cylinder +volume of solid) - Volume of liquid in graduated cylinder                = 5.15-3.45   = 1.7ml

densoty of solid   = mass/volume

                             = 8.436g/1.7ml    = 4.96g/ml >>>>answer